OpenStudy (1645323):

Factor: 2kx + 4mx - 2nx - 3ky - 6my + 3ny Select one: a. (2x - 3)(ky - 2mn + nx) b. (2x - 3y)(k + 2m - n) c. (2x + 3y)(k - 2m + n) d. (2x + 3)(ky + 2mn - nx)

OpenStudy (1645323):

@mathmale @phi

OpenStudy (1645323):

Its B right?

OpenStudy (phi):

I looked at all the choices and just the last term times the last term: a) -3 * nx = -3nx there is no -3nx in the original . So (a) is out b) - 3y * -n= +3ny in the running c) +3y*n= +3ny in the running d) +3*-nx= -3nx out so its b or c

OpenStudy (phi):

b gives the correct -3yk c gives +3yk so b it is. But I did not multiply it out entirely. If there are typos in the question we might not notice.

OpenStudy (1645323):

ok